Student has 10 question test with 5 choices for each answer. He needs to answer 60% or > to pass. If he randomly guesses, what is the % probability he will pass? thank you
Choices are: 0.006 0.377 0.060

Answer by stanbon(75887)

Student has 10 question test with 5 choices for each answer. He needs to answer 60% or > to pass. If he randomly guesses, what is the % probability he will pass? thank you
Choices are: 0.006 0.377 0.060
----
Binomial Problem:
n = 10, p(correct) = 1/5, q(wrong) = 4/5
---
Find P(6 <= x <=10) = 1 - binomcdf(10,1/5,5) = 0.0064
=============
Cheers,
Stan H.